Nine numbers are written in increasing order. The average of the nine numbers is the middle number, the average of the five largest values is 68, and the average of the five smallest values is 44. What is the sum of the nine numbers?
Answer by stanbon(75887)

Nine numbers are written in increasing order. The average of the nine numbers is the middle number, the average of the five largest values is 68, and the average of the five smallest values is 44. What is the sum of the nine numbers?
--------
x-4
x-3
x-2
x-1
x
x+1
x+2
x+3
x+4
------
Middle is average::
Sum/9 = x
9x/9 = x
--------
Average of top 5 = 68
(5x + 10)/5 = 68
5x+10 = 340
5x = 330
x = 66
======
Ans: Sum = 9*66 = 594
========================
Cheers,
Stan H.
========================